# Transfer Learning on Edge Connecting Probability Estimation under Graphon Model

### v1

@@ 5.1 Simulation Experiments
Simulation setup. To evaluate the effectiveness of our proposed method, GTrans, we conduct simulation studies using ten representative graphon functions. Figure 3 illustrates the probability matrices $\mathbf{P}$ of 500 nodes generated for Graphons 1 to 10, where the rows and the columns are ordered by latent position $u$. Graphons 1-5 are graphon functions used in SAS pmlr-v32-chan14 , and graphons 6-10 are used in graphon functions used in ICE qin_iterative_2021 . The details about the structure of these graphon functions can be found in Section E.1.
@@
Impact of density shift. To investigate the performance of GTrans to density differences between source and target domains, we examine scenarios with varying levels of perturbation between source and target graphons: $f_{s}(u,v)=f_{t}(u,v)+\xi$ where $\xi\sim U(0,\lambda)$ if $\lambda>0$, $\xi\sim U(\lambda,0)$ if $\lambda<0$. We vary $\lambda$ from $\{-0.5,-0.45,\ldots,0.5\}$. we truncate the probability if it is larger than 1 or smaller than 0. This allows us to simulate scenarios where the target network is either sparser ($\lambda>0$) or denser ($\lambda<0$) than the source. Figure 6 presents MSE results for two representative graphons (ID 2 and ID 6). Additional results can be found in Figure C1.
@@
For the dense Graphon 2 (density = 0.5), GTrans-GW and GTrans-EGW both show a symmetric U-shaped pattern centered almost around zero, indicating that performance degrades equally whether transferring from a denser or sparser source. In contrast, Graphon 6 is a sparse network (density = 0.13). Here, the performance curve of GTrans-GW and GTrans-EGW achieve its lowest error at a positive value of $\lambda$. This indicates that when the target graph is dense, the impact of density shift on transfer learning is symmetric around zero. However, when the target graph is sparse, transfer performance becomes more sensitive to the direction of the density shift, and transferring from a slightly denser source network yields better results. The underlying reason might be: (1) a denser source network typically exhibits a higher signal-to-noise ratio, yielding a higher-quality initial graphon estimator. (2) When the source network is sparse, $\hat{\mathbf{P}}_{s}^{ini}$ will also be sparse—potentially to an extreme degree where it might even be (approximately) a zero matrix. This severely limits the information available for transfer, regardless of how well the transport plan aligns the domains.
@@
Selection of Hyperparameters. GTrans-GW involves one hyperparameter: threshold $\delta$ while GTrans-GW involves two hyperparameters: threshold $\delta$ and regularization parameters $\epsilon$. In this paper, we employed a network cross-validation procedure developed by li_network_2020 for hyperparameter tuning. (1) Selection of $\delta$: Specifically, when implementing GTrans with either GW or EGW, we choose $\delta$ from a candidate set $\{0.1,0.11,0.12,.....,0.5\}$, and use the cross-validation procedure to select the value that minimizes held-out prediction error. Figures 3(a) and 3(b) illustrate the optimal thresholds identified across different simulation scenarios for GW and EGW, respectively. As we can see: most GW scenarios work best when $\delta=0.15$, while most EGW scenarios work best when $\delta=0.18$. While cross-validation is preferred, we recommend default values of $\delta=0.15$ for GW and $\delta=0.18$ for EGW when tuning is computationally infeasible in practice. (2) Selection of $\epsilon$: When implementing EGW, we select the optimal regularization parameter from candidates $\{0.001,0.005,0.01,\ldots,0.1\}$. Figures C4 illustrates the MSE comparison across different regularization parameters $\epsilon$, highlighting that $\epsilon=0.01$ is consistently effective for most graphons.
@@ 5.2 Application to Real Data
Evaluating probability matrix estimation methods on real networks directly is difficult, since the true probability matrix is unknown. We assess the practical utility of our method by applying it to two downstream applications: graph classification and link prediction. Due to page limitations, we only show results for graph classification. Link prediction results are provided in Appendix D.2.
@@
Results. We adopt the same Graph Convolutional Network (GCN) architecture as used in han2022g , using the same hyperparameters and training procedures for all benchmark comparisons. Full implementation details are provided in Appendix D.1. For each target dataset, we split the dataset into train/validation/test data by $70\%/10\%/20\%$. We report the test accuracy on ten runs. Tables 2 and 3 report test accuracies for our proposed GTrans method against graphon-estimation baselines—classical (NS, USVT, SAS, ICE), resolution-free (GWB xu2021learning , IGNR xia2023implicit , SIGL azizpour2024scalable )—and a graphlet-based method prvzulj2007biological ; shervashidze2009efficient . For GWB/IGNR/SIGL we follow the original implementations and keep hyperparameters unchanged. The graphlet baseline is implemented via GraKeL’s Graphlet Kernel with an SVM on the precomputed kernel matrix; we use graphlet sampling with 700 subgraphs per graph for scalability. We also include two transfer baselines. (1) Pooled Estimation – pooling source & target graphs and estimating a graphon using the G-Mixup procedure with neighborhood smoothing (NS), and (2) Pooled-then-Transfer – using the pooled estimator as the initial source estimate $\hat{P}^{\text{ini}}_{s}$ in our transfer framework, followed by the GTRANS procedure (denoted PGTRANS-GW & PGTRANS-EGW). On IMDB-Binary GTrans-GW achieves 76.30% (transfer from Reddit-Binary) and 76.25% (transfer from COLLAB), while GTrans-EGW further improves to 76.80% and 77.50%, respectively, outperforming all baselines by 2–5%. On IMDB-Multi, GTrans-GW with COLLAB reaches 50.47%, and GTrans-EGW achieves 51.27% with Reddit-Binary as the source, demonstrating clear superiority over all other methods. For PROTEINS-Full, transferring from D&D yields 69.33% with GTrans-GW and 68.52% with GTrans-EGW, consistently outperforming baseline methods by 3–6 points. These results confirm that both GTrans-GW and GTrans-EGW effectively boost classification accuracy by leveraging structural knowledge from larger networks. In particular, from Table 3, both pooled baselines underperform our method, suggesting that naive joint estimation is suboptimal when source and target are misaligned, while our alignment and debiasing steps enable more accurate transfer.
@@
Source Target GTrans-GW GTrans-EGW NS USVT SAS ICE GWB IGNR SIGL Graphlet
Reddit-B IMDB-B 76.30 $\pm$ 2.35 76.80 $\pm$ 1.52 72.90 $\pm$ 2.10 73.85 $\pm$ 2.40 74.25 $\pm$ 1.93 74.30 $\pm$ 2.16 75.30 $\pm$ 2.19 74.35 $\pm$ 3.51 73.50 $\pm$ 2.62 61.10 $\pm$ 2.23
COLLAB IMDB-B 76.25 $\pm$ 2.06 77.50 $\pm$ 2.13 72.90 $\pm$ 2.10 73.85 $\pm$ 2.40 74.25 $\pm$ 1.93 74.30 $\pm$ 2.16 75.30 $\pm$ 2.19 74.35 $\pm$ 3.51 73.50 $\pm$ 2.62 61.10 $\pm$ 2.23
Reddit-B IMDB-M 49.10 $\pm$ 1.33 51.27 $\pm$ 1.98 43.80 $\pm$ 2.59 48.00 $\pm$ 2.93 44.10 $\pm$ 2.05 43.90 $\pm$ 1.27 47.70 $\pm$ 2.53 47.50 $\pm$ 3.32 49.13 $\pm$ 3.04 39.37 $\pm$ 1.62
COLLAB IMDB-M 50.47 $\pm$ 1.42 50.23 $\pm$ 0.92 43.80 $\pm$ 2.59 48.00 $\pm$ 2.93 44.10 $\pm$ 2.05 43.90 $\pm$ 1.27 47.70 $\pm$ 2.53 47.50 $\pm$ 3.32 49.13 $\pm$ 3.04 39.37 $\pm$ 1.62
D&D Proteins 69.33 $\pm$ 2.55 68.52 $\pm$ 1.59 63.18 $\pm$ 1.94 65.11 $\pm$ 2.21 65.25 $\pm$ 1.85 65.38 $\pm$ 1.84 63.45 $\pm$ 2.96 65.87 $\pm$ 2.56 67.13 $\pm$ 2.34 70.11 $\pm$ 2.12
Table 3: Graph classification accuracy (%) comparison between GTRANS and pooled transfer estimation methods (mean $\pm$ std).
@@ 6 Conclusion and Future Work
In this paper, we proposed GTrans, a novel transfer learning framework to estimate the edge connecting probability, addressing the challenge of limited data availability in the target network. By leveraging Gromov-Wasserstein optimal transport, our method aligns latent node structures across source and target domains and effectively transfers graphon information to improve estimation accuracy. Extensive experiments on both synthetic and real-world networks demonstrate that our approach consistently outperforms existing baselines, particularly in scenarios involving small, sparse graphs. By improving estimation accuracy, our framework facilitates various downstream tasks, including data augmentation and link prediction. Despite these promising results, several limitations remain. First, our current framework assumes a single source graph. Extending it to incorporate multiple source networks could further enhance robustness by leveraging diverse structural priors. Second, GTrans is designed for static graphs; future extensions to dynamic or multi-layer networks would enable modeling of time-evolving or multi-modal dependencies. Third, our current formulation does not incorporate node-level covariates. Integrating such covariates could provide valuable auxiliary information for alignment, particularly in domains where topological structure alone may be insufficient for effective transfer. Beyond these technical directions, our framework also opens opportunities for applications across diverse domains. For instance, in systems biology, large gene regulatory or protein interaction networks (e.g., from STRING or BioGRID) can serve as sources for smaller subnetworks in specific tissues or cell types. In online social platforms, massive public graphs (from Facebook, Twitter, or Reddit) could be leveraged to inform smaller-scale communities. In neuroscience, large-scale brain connectomes may transfer knowledge to regional subnetworks. In transportation research, traffic networks of large metropolitan areas (e.g., Chicago, New York) can support modeling in smaller urban settings. These directions can further enhance the flexibility and generalizability of transfer-based graphon estimation.

### v2

@@ 5.1 Simulation Experiments
Simulation setup. To evaluate the effectiveness of our proposed method, GTrans, we conduct simulation studies using ten representative graphon functions. Figure 3 illustrates the probability matrices $\mathbf{P}$ of 500 nodes generated for Graphons 1 to 10, where the rows and the columns are ordered by latent position $u$. Graphons 1-5 are graphon functions used in SAS pmlr-v32-chan14 , and graphons 6-10 are used in graphon functions used in ICE qin_iterative_2021 . The details about the structure of these graphon functions can be found in Section G.1.
@@
Impact of density shift. To investigate the performance of GTrans to density differences
@@
between source and target domains, we examine scenarios with varying levels of perturbation between source and target graphons: $f_{s}(u,v)=f_{t}(u,v)+\xi$ where $\xi\sim U(0,\lambda)$ if $\lambda>0$, $\xi\sim U(\lambda,0)$ if $\lambda<0$. We vary $\lambda$ from $\{-0.5,-0.45,\ldots,0.5\}$, and truncate the probability if it is larger than 1 or smaller than 0. This allows us to simulate scenarios where the target network is either sparser ($\lambda>0$) or denser ($\lambda<0$) than the source. Figure 6 presents MSE results for two representative graphons (ID 2 and ID 6). Additional results can be found in Figure E1. For the dense Graphon 2 (density = 0.5), GTrans-GW and GTrans-EGW both show a symmetric U-shaped pattern centered almost around zero, indicating that performance degrades equally whether transferring from a denser or sparser source. In contrast, Graphon 6 is a sparse network (density = 0.13). Here, the performance curve of GTrans-GW and GTrans-EGW achieve its lowest error at a positive value of $\lambda$. This indicates that when the target graph is dense, the impact of density shift on transfer learning is symmetric around zero. However, when the target graph is sparse, transfer performance becomes more sensitive to the direction of the density shift, and transferring from a slightly denser source network yields better results. The underlying reason might be: (1) a denser source network typically exhibits a higher signal-to-noise ratio, yielding a higher-quality initial graphon estimator.
(2) When the source network is sparse, $\hat{\mathbf{P}}_{s}^{ini}$ will also be sparse—potentially to an extreme degree where it might even be (approximately) a zero matrix. This severely limits the information available for transfer, regardless of how well the transport plan aligns the domains.
@@
Selection of Hyperparameters. GTrans-GW involves one hyperparameter: threshold $\delta$ while GTrans-GW involves two hyperparameters: threshold $\delta$ and regularization parameters $\epsilon$. In this paper, we employed a network cross-validation procedure developed by li_network_2020 for hyperparameter tuning. (1) Selection of $\delta$: Specifically, when implementing GTrans with either GW or EGW, we choose $\delta$ from a candidate set $\{0.1,0.11,0.12,.....,0.5\}$, and use the cross-validation procedure to select the value that minimizes held-out prediction error. Figures 3(a) and 3(b) illustrate the optimal thresholds identified across different simulation scenarios for GW and EGW, respectively. As we can see: most GW scenarios work best when $\delta=0.15$, while most EGW scenarios work best when $\delta=0.18$. While cross-validation is preferred, we recommend default values of $\delta=0.15$ for GW and $\delta=0.18$ for EGW when tuning is computationally infeasible in practice. (2) Selection of $\epsilon$: When implementing EGW, we select the optimal regularization parameter from candidates $\{0.001,0.005,0.01,\ldots,0.1\}$. Figures E4 illustrates the MSE comparison across different regularization parameters $\epsilon$, highlighting that $\epsilon=0.01$ is consistently effective for most graphons.
@@ 5.2 Application to Real Data
Evaluating probability matrix estimation methods on real networks directly is difficult, since the true probability matrix is unknown. We assess the practical utility of our method by applying it to two downstream applications: graph classification and link prediction. Due to page limitations, we only show results for graph classification. Link prediction results are provided in Appendix F.2.
@@
Results. We adopt the same Graph Convolutional Network (GCN) architecture as used in han2022g , using the same hyperparameters and training procedures for all benchmark comparisons. Full implementation details are provided in Appendix F.1. For each target dataset, we split the dataset into train/validation/test data by $70\%/10\%/20\%$. We report the test accuracy on ten runs. Tables 2 report test accuracies for our proposed GTrans method against baselines, including NS  zhang2017estimating , USVT  chatterjee2015matrix , SAS pmlr-v32-chan14 , ICE  qin_iterative_2021 , GWB xu2021learning , IGNR xia2023implicit , SIGL azizpour2024scalable , and graphlet prvzulj2007biological ; shervashidze2009efficient . For GWB/IGNR/SIGL we follow the original implementations and keep hyperparameters unchanged. The graphlet baseline is implemented via GraKeL’s Graphlet Kernel with an SVM on the precomputed kernel matrix; we use graphlet sampling with 700 subgraphs per graph for scalability. On IMDB-Binary GTrans-GW achieves 76.30% (transfer from Reddit-Binary) and 76.25% (transfer from COLLAB), while GTrans-EGW further improves to 76.80% and 77.50%, respectively, outperforming all baselines by 2–5%. On IMDB-Multi, GTrans-GW with COLLAB reaches 50.47%, and GTrans-EGW achieves 51.27% with Reddit-Binary as the source, demonstrating clear superiority over all other methods. For PROTEINS-Full, transferring from D&D yields 69.33% with GTrans-GW and 68.52% with GTrans-EGW, consistently outperforming baseline methods by 3–6 points. These results confirm that both GTrans-GW and GTrans-EGW effectively boost classification accuracy by leveraging structural knowledge from larger networks.
@@
Source Target GTrans-GW GTrans-EGW NS USVT SAS
Reddit-B IMDB-B 76.30 $\pm$ 2.35 76.80 $\pm$ 1.52 72.90 $\pm$ 2.10 73.85 $\pm$ 2.40 74.25 $\pm$ 1.93
COLLAB IMDB-B 76.25 $\pm$ 2.06 77.50 $\pm$ 2.13 72.90 $\pm$ 2.10 73.85 $\pm$ 2.40 74.25 $\pm$ 1.93
Reddit-B IMDB-M 49.10 $\pm$ 1.33 51.27 $\pm$ 1.98 43.80 $\pm$ 2.59 48.00 $\pm$ 2.93 44.10 $\pm$ 2.05
COLLAB IMDB-M 50.47 $\pm$ 1.42 50.23 $\pm$ 0.92 43.80 $\pm$ 2.59 48.00 $\pm$ 2.93 44.10 $\pm$ 2.05
D&D Proteins 69.33 $\pm$ 2.55 68.52 $\pm$ 1.59 63.18 $\pm$ 1.94 65.11 $\pm$ 2.21 65.25 $\pm$ 1.85
Source Target ICE GWB IGNR SIGL Graphlet
Reddit-B IMDB-B 74.30 $\pm$ 2.16 75.30 $\pm$ 2.19 74.35 $\pm$ 3.51 73.50 $\pm$ 2.62 61.10 $\pm$ 2.23
COLLAB IMDB-B 74.30 $\pm$ 2.16 75.30 $\pm$ 2.19 74.35 $\pm$ 3.51 73.50 $\pm$ 2.62 61.10 $\pm$ 2.23
Reddit-B IMDB-M 43.90 $\pm$ 1.27 47.70 $\pm$ 2.53 47.50 $\pm$ 3.32 49.13 $\pm$ 3.04 39.37 $\pm$ 1.62
COLLAB IMDB-M 43.90 $\pm$ 1.27 47.70 $\pm$ 2.53 47.50 $\pm$ 3.32 49.13 $\pm$ 3.04 39.37 $\pm$ 1.62
D&D Proteins 65.38 $\pm$ 1.84 63.45 $\pm$ 2.96 65.87 $\pm$ 2.56 67.13 $\pm$ 2.34 70.11 $\pm$ 2.12
@@ 6 Conclusion and Future Work
In this paper, we proposed GTrans, a novel transfer learning framework to estimate the edge connecting probability, addressing the challenge of limited data availability in the target network. By leveraging Gromov-Wasserstein optimal transport, our method aligns latent node structures across source and target domains and effectively transfers graphon information to improve estimation accuracy. Extensive experiments on both synthetic and real-world networks demonstrate that our approach consistently outperforms existing baselines, particularly in scenarios involving small, sparse graphs. By improving estimation accuracy, our framework facilitates various downstream tasks, including data augmentation and link prediction. Despite these promising results, several limitations remain. First, our current framework assumes a single source graph. Extending it to incorporate multiple source networks could further enhance robustness by leveraging diverse structural priors. Second, GTrans is designed for static graphs; future extensions to dynamic or multi-layer networks would enable modeling of time-evolving or multi-modal dependencies. Third, our current formulation does not incorporate node-level covariates. Integrating such covariates could provide valuable auxiliary information for alignment, particularly in domains where topological structure alone may be insufficient for effective transfer. These directions can further enhance the flexibility and generalizability of transfer-based graphon estimation.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The scope and contributions of this work were included in the abstract and introduction, particularly highlighted at the end of the Introduction section.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: Limitations and future directions of the work were included in the Discussion section.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: Full assumptions and proof were included in the Theoretical Properties section and the Appendix.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: The proposed algorithms and implementation details of experiments were described in the Experiments Section and Appendix.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: The real data used in this work are publicly accessible and properly cited. We also include the code and descriptions in the supplementary materials.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: The details were discussed in the Experiments section and Appendix.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: In the Experiments section and Appendix, the error bars, standard deviation, and number of replicates, were reported and clearly stated.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We provide detailed descriptions of the computational resources used for each experiment in the Appendix. Furthermore, our open-sourced code includes configuration files and parameter settings to replicate the experiments efficiently.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The work of this paper is conducted with NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: Societal impacts of the work were included in the Introduction section.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: The paper poses no such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: Existing assets were cited and credited throughout this paper.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: Documentation was provided with the supplementary code.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: This paper does not conduct research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: This paper does not conduct research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: The core methods presented in this research do not utilize Large LLMs as an essential, original, or non-standard component. LLMs were not used in any part of the model design, experimental evaluation, or optimization processes. Additionally, LLMs were not employed for data augmentation or pre-processing in our experiments.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.